# Are Lithium-Rich Giants Binaries?: A Radial Velocity Variability Analysis of 1,400 Giants

### v2

@@ 5 Summary and conclusions
We develop a method for detecting radial velocity (RV) variability when using RVs from heterogeneous sources, to apply it to low-mass field giants with well-determined lithium (Li) content and explore the question of whether the phenomenon of Li-rich giants is a result of binary stellar evolution. The method is a variation of the F2 statistic and works by comparing three independent RVs from different epochs, accounting for the corresponding uncertainties. We test and calibrate the sensitivity of the method against two control samples, a catalog of RV standard stars and another of confirmed spectroscopic binaries, to evaluate the degree of confidence in identifying single (RV non-variable) and binary (RV variable) stars. We find, non-surprisingly, that the accuracy of the method in correctly separating the two control samples mainly depends on the precision of the RVs available for the single stars. The separation achieves a 100% success rate when only high-precision RVs are used for the RV standards (the ”single” stars), and diminishes to 80-85% when the less-precise RVs from RAVE and Gaia are included in the RV variability analysis (see Figures 2 and 3). Since these are the data that we will use for the analysis of giants with Li content, this is the level of certainty that our ability to identify binaries among our sample of giants will have. We also compare our results with the information on binarity and variability available in Gaia DR3, but find that the available information is insufficient for our purposes because the survey has so far flagged only a small fraction of binary systems.
We assemble all stars with RV determinations from Gaia DR3, RAVE DR6, and GALAH DR3, and, following the recipe from the GALAH team itself, we use stellar parameters from the latter to make a selection of 1418 giants with good Li determinations. For 1030 of these we can determine evolutionary states, producing subsamples of 563 first ascent red giants (RGBs) and 467 red clump (RC) giants (see Figure 4). This is the largest sample of low-mass giants with well determined Li content ever studied for binarity purposes. We apply our RV variability classification method to these samples, and our main findings can be summarized as follows:
@@
    the fraction of giants (RGB+RC) with high certainty of RV variability is $\sim$33%, which is of the order of the close binary fraction observed for main-sequence stars (e.g. Gao et al. 2017; Moe et al. 2019) and giants (Badenes et al. 2018) at typical disk metallicities. This provides evidence, additional to that obtained from the control samples, that our classification analysis is reliable.
@@
    the fraction of RV variable stars in the full sample (RGB+RC; left-most panel in Figure 5) is about the same for Li-rich giants (A(Li) $>$ 1.5) and for Li-normal giants (A(Li) $<$ 1.5). Moreover, there is no excess of RV variable Li-rich giants with extreme Li content (the super Li-rich giants, with A(Li) $>$ 2.7 or A(Li) $>$ 3.3). All this argues against a binary interaction phenomenon as the main mechanism for explaining Li-rich giants in general.
@@
    the fraction of Li-rich giants is more than two times larger in the RC than in the RGB. Although we do not attempt to quantify selection effects in our samples with respect to Li content, we note that this is consistent with the fact that most Li-rich giants with seismic based determinations of mass and evolutionary state have turned out to be clump giants (e.g. Silva Aguirre et al. 2014; Deepak & Reddy 2019; Deepak & Lambert 2021; Mallick et al. 2023). Still, the incidence of Li-rich objects among 1st-ascent RGB giants in our sample ($\sim$ 11%) is not small, and they too remain unexplained by standard stellar evolution.
@@
    when focusing on the RC population (right-most panel in Figure 5), we find that the fractions of Li-rich and of Li-normal RC giants are the same across all RV variability classes, again likely indicating that the close binary fraction of Li-rich giants in the RC is not special. This is a highly significant result statistically speaking, obtained for the largest sample of RC giants with well-determined Li content ever analyzed for these purposes, and keeps adding to the evidence against a binary-related genesis for Li-rich giants in the RC. For super Li-rich RC giants, the statistics become poorer, but we observe that the fraction of highly RV-variable objects of that type appears to decrease rapidly with increasing Li content.
@@
    while the binary fraction in the RC does not appear to depend on Li content, we find evidence that this may not be the case for 1st ascent RGB giants, for which we obtain that about half of the Li-rich and super Li-rich objects in this evolutionary state have high RV variability (middle panel in Figure 5). This is an intriguing result that contrasts with what we find for the RC, and which may be signaling some causal connection between binarity and Li production during the first-ascent RGB. However, as there are only half as many RGB giants as RC giants in our sample, this is a result not as statistically significant as that of the RC. Future work with a larger number of RV epochs and a cadence better suited for binary detection and characterization would be needed to confirm or rule out this specific result with a higher degree of confidence.
Thus the main conclusion of the present work needs to be separated by evolutionary state. In the case of Li-rich giants in the clump, our RV variability analysis argues against binaries having to do with the origin of the excess Li. This lends support to work that provides indirect but compelling evidence for a causal relation with the core helium flash (Mallick et al. 2023; del Moral & Chanamé 2024, in preparation). In the case of Li-rich giants on the 1st-ascent RGB, instead, our analysis leaves some open room for a binary-related origin of the excess Li, but more RV measurements are needed for more robust conclusions for that evolutionary state. We stress that this apparent dichotomy between Li-rich objects in the RC and the RGB should not be seen necessarily as a problem, because, as discussed in Chanamé et al. (2022), today’s field red giants are different than the progenitors of today’s field red clump giants. Field clump giants are a younger population, with a stellar mass distribution that extends to higher masses than the older field red giants we observe today, and thus it may well be the case that Li-rich giants on the RGB have a different origin than Li-rich giants in the clump.
Acknowledgements
M.C.-T. acknowledges support from the Trottier Space Institute (TSI) and the Centre de Recherche en Astrophysique du Québec (CRAQ). J.C. acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) via Proyecto Fondecyt Regular 1231345; and from ANID BASAL projects CATA-Puente ACE210002 and CATA2-FB210003.

### v3

@@ 5 Summary and conclusions
We developed a method for detecting RV variability when using RVs from heterogeneous sources, to apply it to low-mass field giants with a well-determined Li content and explored the question of whether the phenomenon of Li-rich giants is a result of binary stellar evolution. The method is a variation of the F2 statistic and works by comparing three independent RVs from different epochs, accounting for the corresponding uncertainties. We tested and calibrated the sensitivity of the method against two control samples, a catalog of RV standard stars and another of confirmed spectroscopic binaries, to evaluate the degree of confidence in identifying single (RV non-variable) and binary (RV variable) stars. We find, non-surprisingly, that the accuracy of the method in correctly separating the two control samples mainly depends on the precision of the RVs available for the single stars. The separation achieves a 100% success rate when only high-precision RVs are used for the RV standards (the ”single” stars), and diminishes to 80-85% when the less precise RVs from RAVE and Gaia are included in the RV variability analysis (see Figs. 2 and 3). Since these are the data that we will use for the analysis of giants with a Li content, this is the level of certainty that our ability to identify binaries among our sample of giants will have. We also compared our results with the information on binarity and variability available in Gaia DR3, but find that the available information is insufficient for our purposes because the survey has thus far only flagged a small fraction of binary systems.
We assembled all stars with RV determinations from Gaia DR3, RAVE DR6, and GALAH DR3, and, following the recipe from the GALAH team itself, we used stellar parameters from the latter to make a selection of 1418 giants with good Li determinations. For 1030 of these, we could determine evolutionary states, producing subsamples of 563 first ascent red giants (RGBs) and 467 red clump giants (see Fig. 4). This is the largest sample of low-mass giants with a well-determined Li content ever studied for binarity purposes. We applied our RV variability classification method to these samples, and our main findings are summarized hereafter.
@@
    The fraction of giants (RGB+RC) with a high certainty of RV variability is $\sim$33%, which is on the order of the close binary fraction observed for main-sequence stars (e.g., Gao et al. 2017; Moe et al. 2019) and giants (Badenes et al. 2018) at typical disk metallicities. This provides evidence, in addition to that obtained from the control samples, that our classification analysis is reliable.
@@
    The fraction of RV variable stars in the full sample (RGB+RC; left-most panel in Fig. 5) is about the same for Li-rich giants (A(Li) $>$ 1.5) and for Li-normal giants (A(Li) $<$ 1.5). Moreover, there is no excess of RV variable Li-rich giants with an extreme Li content (the super Li-rich giants, with A(Li) $>$ 2.7 or A(Li) $>$ 3.3). All this argues against a binary interaction phenomenon as the main mechanism for explaining Li-rich giants in general.
@@
    The fraction of Li-rich giants is more than two times larger in the RC than in the RGB. Although we did not attempt to quantify selection effects in our samples with respect to the Li content, we note that this is consistent with the fact that most Li-rich giants with seismic-based determinations of mass and evolutionary state have turned out to be clump giants (e.g., Silva Aguirre et al. 2014; Deepak & Reddy 2019; Deepak & Lambert 2021; Mallick et al. 2023). Even so, the incidence of Li-rich objects among first-ascent RGB giants in our sample ($\sim$ 11%) is not small, and they too remain unexplained by standard stellar evolution.
@@
    When focusing on the RC population (right-most panel in Fig. 5), we find that the fractions of Li-rich and of Li-normal RC giants are the same across all RV variability classes, again likely indicating that the close binary fraction of Li-rich giants in the RC is not special. This is a highly significant result statistically speaking, obtained for the largest sample of RC giants with a well-determined Li content ever analyzed for these purposes, and this keeps adding to the evidence against a binary-related genesis for Li-rich giants in the RC. For super Li-rich RC giants, the statistics become poorer, but we observe that the fraction of highly RV-variable objects of that type appears to decrease rapidly with an increasing Li content.
@@
    While the binary fraction in the RC does not appear to depend on the Li content, we find evidence that this may not be the case for first-ascent RGB giants, for which we determine that about half of the Li-rich and super Li-rich objects in this evolutionary state have a high RV variability (middle panel in Fig. 5). This is an intriguing result that contrasts with what we find for the RC, and which may be signaling some causal connection between binarity and Li production during the first-ascent RGB. However, as there are only half as many RGB giants as RC giants in our sample, this is a result not as statistically significant as that of the RC. Future work with a larger number of RV epochs and a cadence better suited for binary detection and characterization would be needed to confirm or rule out this specific result with a higher degree of confidence.
Thus the main conclusion of the present work needs to be separated by the evolutionary state. In the case of Li-rich giants in the clump, our RV variability analysis argues against binaries having to do with the origin of the excess Li. This lends support to work that provides indirect but compelling evidence for a causal relation with the core helium flash (Mallick et al. 2023; del Moral & Chanamé, in prep.). In the case of Li-rich giants on the first-ascent RGB, instead, our analysis leaves some open room for a binary-related origin of the excess Li, but more RV measurements are needed for more robust conclusions for that evolutionary state. We stress that this apparent dichotomy between Li-rich objects in the RC and the RGB should not necessarily be seen as a problem, because, as discussed in Chanamé et al. (2022), today’s field red giants are different than the progenitors of today’s field red clump giants. Field clump giants are a younger population, with a stellar mass distribution that extends to higher masses than the older field red giants we observe today, and thus it may well be the case that Li-rich giants on the RGB have a different origin than Li-rich giants in the clump.
Acknowledgements.
M.C.-T. acknowledges support from the Trottier Space Institute (TSI) and the Centre de Recherche en Astrophysique du Québec (CRAQ). J.C. acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) via Proyecto Fondecyt Regular 1231345; and from ANID BASAL projects CATA-Puente ACE210002 and CATA2-FB210003. We thank the anonymous referee for the insightful suggestions on the previous versions of the paper.